A mechanic has a supply of 25% antifreeze solution and 50% antifreeze solution. How much of each should be mixed together in order to have 20 gallons of a 35% antifreeze solution
Answer by richwmiller(17219)

a+b=20
0.25*a+0.5*b=0.35*20
a=12., b=8.